# Chronos: Learning the Language of Time Series

### v2

@@ Appendix E Additional Results
Australian Electricity 0.068 0.079 0.077 0.069 0.082 0.069 0.097 0.054 0.046 0.037 0.087 0.052 0.036 0.066 0.034 0.038 0.125 0.055 0.073 0.084 0.159
Car Parts 1.067 1.058 1.033 1.029 1.031 - 1.011 1.652 1.626 0.998 0.967 0.941 0.871 1.119 0.880 0.877 1.309 1.337 - 1.600 -
CIF 2016 0.013 0.013 0.014 0.015 0.014 0.014 0.041 0.011 0.013 0.140 0.136 0.086 0.011 0.033 0.032 0.039 0.039 0.027 0.017 0.015 0.009
Covid Deaths 0.045 0.046 0.060 0.078 0.089 0.032 0.276 0.038 0.034 0.065 0.108 0.918 0.034 0.077 0.038 0.056 0.064 0.094 0.029 0.133 0.133
Dominick 0.337 0.338 0.344 0.351 0.341 - 0.443 0.361 0.345 0.345 0.364 0.327 0.320 0.435 0.313 0.312 0.483 0.485 - 0.453 0.453
ERCOT Load 0.016 0.015 0.017 0.018 0.016 0.053 0.033 0.019 0.021 0.017 0.032 0.024 0.023 0.023 0.020 0.020 0.122 0.041 0.052 0.037 0.181
ETT (15 Min.) 0.064 0.065 0.060 0.065 0.068 0.088 0.080 0.074 0.070 0.054 0.069 0.113 0.075 0.071 0.051 0.053 0.095 0.079 0.073 0.141 0.121
ETT (Hourly) 0.074 0.075 0.077 0.083 0.076 0.122 0.106 0.094 0.083 0.071 0.081 0.142 0.082 0.076 0.081 0.074 0.132 0.133 0.105 0.122 0.202
Exchange Rate 0.017 0.019 0.020 0.017 0.017 0.015 0.011 0.010 0.012 0.010 0.009 0.016 0.011 0.008 0.010 0.011 0.010 0.010 0.011 0.013 0.015
FRED-MD 0.026 0.024 0.017 0.019 0.029 0.041 0.389 0.050 0.046 0.042 0.043 0.058 0.112 0.069 0.057 0.061 0.055 0.057 0.056 0.122 0.064
Hospital 0.057 0.057 0.058 0.059 0.059 0.066 0.093 0.059 0.057 0.070 0.056 0.064 0.053 0.089 0.052 0.050 0.053 0.055 0.058 0.073 0.087
M1 (Monthly) 0.129 0.126 0.138 0.141 0.135 0.181 0.196 0.155 0.149 0.165 0.150 0.150 0.175 0.189 0.189 0.187 0.162 0.159 0.146 0.191 0.258
M1 (Quarterly) 0.105 0.099 0.103 0.101 0.115 0.115 0.141 0.106 0.106 0.078 0.089 0.094 0.122 0.079 0.111 0.085 0.083 0.082 0.091 0.150 0.130
M1 (Yearly) 0.176 0.183 0.169 0.179 0.208 0.144 0.293 0.195 0.207 0.165 0.139 0.168 0.124 0.245 0.198 0.182 0.142 0.137 0.160 0.209 0.209
M3 (Monthly) 0.096 0.097 0.099 0.099 0.104 0.108 0.155 0.102 0.101 0.113 0.099 0.100 0.096 0.121 0.097 0.101 0.093 0.095 0.102 0.149 0.158
M3 (Quarterly) 0.073 0.075 0.078 0.080 0.078 0.084 0.134 0.080 0.084 0.074 0.073 0.072 0.071 0.086 0.076 0.080 0.069 0.070 0.079 0.101 0.103
M3 (Yearly) 0.150 0.151 0.153 0.158 0.149 0.148 0.192 0.165 0.170 0.133 0.122 0.130 0.130 0.143 0.182 0.181 0.127 0.128 0.162 0.167 0.167
M4 (Quarterly) 0.082 0.083 0.083 0.085 0.086 - 0.132 0.081 0.080 0.074 0.080 0.079 0.080 0.085 0.073 0.073 0.080 0.079 0.082 0.119 0.110
M4 (Yearly) 0.133 0.135 0.135 0.139 0.147 - 0.178 0.121 0.139 0.106 0.111 0.109 0.110 0.115 - - 0.118 0.115 0.130 0.161 0.161
M5 0.588 0.587 0.591 0.596 0.600 - 0.635 0.692 0.584 0.597 0.657 0.594 0.560 0.687 0.563 0.560 0.628 0.636 0.624 1.024 1.024
NN5 (Daily) 0.155 0.160 0.170 0.172 0.165 0.242 0.261 0.181 0.162 0.149 0.155 0.154 0.145 0.159 0.149 0.147 0.264 0.294 0.312 0.425 0.425
NN5 (Weekly) 0.089 0.090 0.090 0.089 0.094 0.092 0.111 0.092 0.092 0.081 0.087 0.098 0.086 0.090 0.098 0.114 0.088 0.090 0.090 0.123 0.123
Tourism (Monthly) 0.099 0.099 0.112 0.108 0.096 0.125 0.213 0.125 0.114 0.092 0.092 0.104 0.096 0.101 0.092 0.084 0.090 0.091 0.093 0.104 0.297
Tourism (Quarterly) 0.062 0.068 0.070 0.077 0.068 0.071 0.202 0.099 0.085 0.074 0.072 0.082 0.074 0.080 0.077 0.063 0.070 0.061 0.098 0.119 0.166
Tourism (Yearly) 0.185 0.199 0.197 0.217 0.187 0.163 0.238 0.167 0.164 0.136 0.127 0.179 0.102 0.165 0.139 0.154 0.159 0.176 0.156 0.209 0.209
Traffic 0.254 0.261 0.261 0.262 0.252 0.287 0.256 0.225 0.232 0.246 0.233 0.234 0.264 0.250 0.263 0.270 0.557 0.905 - 0.362 0.643
Weather 0.139 0.140 0.143 0.149 0.145 - 0.164 0.135 0.132 0.143 0.147 0.152 0.151 0.174 0.143 0.144 0.214 0.217 0.185 0.217 0.217
Agg. Relative Score 0.649 0.661 0.672 0.690 0.700 0.804 1.097 0.699 0.685 0.684 0.733 0.842 0.639 0.757 0.672 0.681 0.838 0.793 0.761 1.000 1.152
Avg. Rank 7.296 8.519 9.630 10.926 10.815 14.426 17.444 11.222 10.333 6.815 8.074 10.704 6.778 12.148 8.667 8.333 10.370 10.037 12.204 17.815 18.444
@@
Australian Electricity 1.306 1.333 1.403 1.212 1.370 1.186 2.158 1.635 1.241 0.981 0.871 1.473 0.997 0.810 1.278 0.794 0.828 1.161 2.391 0.897 1.393 1.253 2.362
Car Parts 0.911 0.897 0.891 0.893 0.879 - 2.657 0.816 1.734 1.538 0.803 0.798 0.817 0.799 0.879 0.803 0.803 0.891 1.185 1.229 - 1.201 -
CIF 2016 0.985 0.995 1.016 1.052 1.065 1.384 3.588 2.235 1.218 1.115 1.537 1.363 1.309 1.553 1.145 1.389 1.440 0.960 0.957 1.002 1.006 1.289 1.263
Covid Deaths 42.762 42.641 42.689 43.525 48.028 32.143 91.515 78.456 33.010 33.019 36.465 38.203 102.457 30.635 40.418 31.771 31.730 75.909 38.114 45.407 31.705 46.912 46.912
Dominick 0.837 0.838 0.838 0.853 0.841 - 3.274 1.250 0.880 0.845 0.867 0.851 0.812 0.800 0.880 0.782 0.782 1.813 0.885 1.016 - 0.871 0.871
ERCOT Load 0.556 0.541 0.577 0.579 0.568 1.319 3.975 0.834 0.596 0.653 0.553 1.197 0.780 0.690 0.651 0.615 0.648 0.558 2.826 1.306 1.284 0.761 4.234
ETT (15 Min.) 0.712 0.710 0.669 0.732 0.753 1.042 1.138 0.967 0.956 0.777 0.652 0.874 1.339 0.962 0.724 0.643 0.659 0.574 1.183 0.583 0.879 1.169 1.164
ETT (Hourly) 0.738 0.749 0.769 0.794 0.750 1.232 1.833 1.002 0.887 0.833 0.729 0.814 1.509 0.875 0.695 0.811 0.782 0.768 1.139 0.900 0.977 0.932 1.651
Exchange Rate 3.231 3.460 3.357 3.223 3.206 1.743 7.583 3.087 1.565 1.937 1.540 1.615 3.105 2.361 1.459 2.041 2.149 2.709 1.643 1.648 1.882 1.740 1.874
FRED-MD 0.592 0.584 0.576 0.537 0.569 0.513 2.621 2.283 0.611 0.603 0.745 0.621 0.849 0.929 0.713 0.696 0.635 0.693 0.544 0.566 0.473 1.101 0.622
Hospital 0.815 0.820 0.819 0.821 0.833 0.861 1.775 0.939 0.821 0.823 0.859 0.804 0.857 0.799 0.940 0.781 0.760 0.793 0.760 0.761 0.820 0.921 0.968
M1 (Monthly) 1.086 1.119 1.163 1.172 1.164 1.415 2.172 1.875 1.271 1.243 1.208 1.122 1.266 1.326 1.369 1.333 1.236 1.198 1.072 1.099 1.153 1.314 1.468
M1 (Quarterly) 1.699 1.735 1.776 1.799 1.767 1.802 9.931 3.036 1.858 1.818 1.920 1.741 1.904 2.144 1.943 2.061 2.043 1.958 1.710 1.683 1.770 2.078 1.952
M1 (Yearly) 4.296 4.582 4.616 4.898 4.674 4.077 23.089 7.149 4.635 4.707 4.042 3.685 4.727 4.316 11.565 5.568 6.212 3.675 4.110 3.697 3.870 4.894 4.894
M3 (Monthly) 0.853 0.861 0.883 0.898 0.925 0.996 2.240 1.846 0.948 0.924 1.225 0.943 0.950 0.916 1.161 0.899 0.883 0.950 0.869 0.861 0.933 1.146 1.175
M3 (Quarterly) 1.170 1.185 1.250 1.270 1.230 1.450 10.176 2.886 1.439 1.449 1.264 1.209 1.257 1.160 1.572 1.202 1.147 1.448 1.125 1.130 1.419 1.425 1.464
M3 (Yearly) 3.094 3.186 3.238 3.348 3.112 3.140 18.728 5.114 3.647 3.824 2.949 2.827 3.026 2.860 3.435 3.432 3.547 3.418 2.696 2.613 3.165 3.172 3.172
M4 (Quarterly) 1.203 1.213 1.228 1.252 1.285 - 6.927 2.663 1.284 1.259 1.150 1.254 1.241 1.248 1.229 1.157 1.129 1.215 1.188 1.193 1.276 1.602 1.477
M4 (Yearly) 3.569 3.641 3.613 3.705 3.894 - - 5.866 3.603 4.173 3.072 3.178 3.221 3.119 3.295 - - 3.374 3.374 3.124 3.730 3.974 3.974
M5 0.946 0.942 0.942 0.946 0.972 - 1.530 0.965 1.439 0.930 0.919 0.956 0.959 0.909 1.027 0.917 0.917 0.935 1.101 1.100 1.057 1.399 1.399
NN5 (Daily) 0.570 0.584 0.620 0.638 0.612 0.953 1.375 0.992 0.702 0.627 0.575 0.585 0.585 0.556 0.604 0.571 0.571 0.720 1.039 1.073 1.214 1.292 1.292
NN5 (Weekly) 0.917 0.925 0.930 0.927 0.962 0.968 1.349 1.141 1.003 0.979 0.877 0.920 1.034 0.896 0.966 0.919 1.014 1.268 0.978 0.984 0.995 1.063 1.063
Tourism (Monthly) 1.741 1.817 1.891 1.937 1.772 2.139 4.348 3.030 2.042 1.913 1.572 1.529 1.629 1.686 1.551 1.514 1.486 1.573 1.497 1.680 1.573 1.631 3.591
Tourism (Quarterly) 1.660 1.705 1.727 1.822 1.814 1.916 5.595 3.695 2.702 2.331 1.723 1.586 1.769 1.729 1.690 1.585 1.618 1.750 1.590 1.658 1.661 1.699 3.633
Tourism (Yearly) 3.729 3.858 3.879 4.049 3.839 3.309 12.093 3.755 3.074 3.267 3.138 3.702 4.130 3.047 3.406 3.448 3.564 - 3.138 3.078 4.043 3.552 3.552
Traffic 0.798 0.823 0.833 0.847 0.810 0.973 1.909 0.829 0.725 0.760 0.790 0.737 0.797 0.880 0.821 0.927 0.968 0.787 1.685 1.794 - 1.077 2.052
Weather 0.827 0.829 0.839 0.855 0.871 - 2.003 1.001 0.831 0.808 0.860 0.911 0.945 0.913 0.997 0.910 0.888 0.972 1.079 0.991 0.907 1.004 1.004
Agg. Relative Score 0.831 0.844 0.856 0.866 0.866 0.962 2.450 1.291 0.909 0.874 0.810 0.843 0.951 0.847 0.894 0.830 0.835 0.895 0.953 0.875 0.908 1.000 1.188
Avg. Rank 7.296 8.704 10.259 11.741 11.222 15.630 22.241 18.741 12.593 11.741 7.630 8.556 13.296 9.111 12.259 9.019 9.204 11.259 9.852 8.741 12.648 16.000 18.259

### v3

@@ Appendix E Additional Results
Australian Electricity 0.067 0.075 0.074 0.063 0.078 0.069 0.097 0.055 0.046 0.037 0.087 0.052 0.036 0.066 0.034 0.038 0.070 0.125 0.055 0.073 0.084 0.159
Car Parts 1.060 1.057 1.029 1.024 1.028 - 1.011 1.655 1.617 0.998 0.967 0.941 0.871 1.119 0.880 0.877 1.283 1.309 1.337 - 1.600 -
CIF 2016 0.014 0.013 0.015 0.013 0.015 0.014 0.041 0.010 0.048 0.140 0.136 0.086 0.011 0.033 0.032 0.039 0.024 0.039 0.027 0.017 0.015 0.009
Covid Deaths 0.045 0.048 0.059 0.084 0.079 0.032 0.276 0.038 0.035 0.065 0.108 0.918 0.034 0.077 0.038 0.056 0.037 0.064 0.094 0.029 0.133 0.133
Dominick 0.332 0.333 0.338 0.346 0.336 - 0.443 0.361 0.346 0.345 0.364 0.327 0.320 0.435 0.313 0.312 0.439 0.483 0.485 - 0.453 0.453
ERCOT Load 0.019 0.016 0.018 0.018 0.017 0.053 0.033 0.019 0.022 0.017 0.032 0.024 0.023 0.023 0.020 0.020 0.050 0.122 0.041 0.052 0.037 0.181
ETT (15 Min.) 0.068 0.069 0.064 0.072 0.073 0.088 0.080 0.075 0.069 0.054 0.069 0.113 0.075 0.071 0.051 0.053 0.061 0.095 0.079 0.073 0.141 0.121
ETT (Hourly) 0.073 0.081 0.080 0.085 0.080 0.122 0.106 0.096 0.085 0.071 0.081 0.142 0.082 0.076 0.081 0.074 0.087 0.132 0.133 0.105 0.122 0.202
Exchange Rate 0.013 0.014 0.013 0.012 0.013 0.015 0.011 0.010 0.012 0.010 0.009 0.016 0.011 0.008 0.010 0.011 0.011 0.010 0.010 0.011 0.013 0.015
FRED-MD 0.020 0.022 0.017 0.017 0.022 0.041 0.389 0.045 0.049 0.042 0.043 0.058 0.112 0.069 0.057 0.061 0.059 0.055 0.057 0.056 0.122 0.064
Hospital 0.056 0.056 0.057 0.058 0.057 0.066 0.093 0.060 0.057 0.070 0.056 0.064 0.053 0.089 0.052 0.050 0.052 0.053 0.055 0.058 0.073 0.087
M1 (Monthly) 0.130 0.128 0.139 0.138 0.131 0.181 0.196 0.155 0.154 0.165 0.150 0.150 0.175 0.189 0.189 0.187 0.162 0.162 0.159 0.146 0.191 0.258
M1 (Quarterly) 0.107 0.105 0.103 0.103 0.116 0.115 0.141 0.111 0.107 0.078 0.089 0.094 0.122 0.079 0.111 0.085 0.083 0.083 0.082 0.091 0.150 0.130
M1 (Yearly) 0.183 0.181 0.172 0.179 0.204 0.144 0.293 0.194 0.190 0.165 0.139 0.168 0.124 0.245 0.198 0.182 0.135 0.142 0.137 0.160 0.209 0.209
M3 (Monthly) 0.096 0.097 0.100 0.099 0.106 0.108 0.155 0.102 0.101 0.113 0.099 0.100 0.096 0.121 0.097 0.101 0.094 0.093 0.095 0.102 0.149 0.158
M3 (Quarterly) 0.074 0.076 0.079 0.081 0.078 0.084 0.134 0.080 0.085 0.074 0.073 0.072 0.071 0.086 0.076 0.080 0.072 0.069 0.070 0.079 0.101 0.103
M3 (Yearly) 0.151 0.153 0.155 0.159 0.148 0.148 0.192 0.167 0.170 0.133 0.122 0.130 0.130 0.143 0.182 0.181 0.144 0.127 0.128 0.162 0.167 0.167
M4 (Quarterly) 0.082 0.083 0.084 0.086 0.087 - 0.132 0.081 0.080 0.074 0.080 0.079 0.080 0.085 0.073 0.073 0.079 0.080 0.079 0.082 0.119 0.110
M4 (Yearly) 0.134 0.137 0.136 0.140 0.148 - 0.178 0.121 0.138 0.106 0.111 0.109 0.110 0.115 - - 0.114 0.118 0.115 0.130 0.161 0.161
M5 0.587 0.586 0.590 0.595 0.598 - 0.635 0.692 0.584 0.597 0.657 0.594 0.560 0.687 0.563 0.560 0.653 0.628 0.636 0.624 1.024 1.024
NN5 (Daily) 0.156 0.161 0.169 0.173 0.162 0.242 0.261 0.181 0.162 0.149 0.155 0.154 0.145 0.159 0.149 0.147 0.293 0.264 0.294 0.312 0.425 0.425
NN5 (Weekly) 0.091 0.091 0.090 0.091 0.094 0.092 0.111 0.092 0.093 0.081 0.087 0.098 0.086 0.090 0.098 0.114 0.092 0.088 0.090 0.090 0.123 0.123
Tourism (Monthly) 0.100 0.103 0.113 0.109 0.095 0.125 0.213 0.121 0.111 0.092 0.092 0.104 0.096 0.101 0.092 0.084 0.083 0.090 0.091 0.093 0.104 0.297
Tourism (Quarterly) 0.061 0.069 0.069 0.074 0.068 0.071 0.202 0.100 0.085 0.074 0.072 0.082 0.074 0.080 0.077 0.063 0.075 0.070 0.061 0.098 0.119 0.166
Tourism (Yearly) 0.183 0.207 0.200 0.218 0.194 0.163 0.238 0.168 0.161 0.136 0.127 0.179 0.102 0.165 0.139 0.154 0.162 0.159 0.176 0.156 0.209 0.209
Traffic 0.256 0.264 0.263 0.264 0.254 0.287 0.256 0.225 0.231 0.246 0.233 0.234 0.264 0.250 0.263 0.270 - 0.557 0.905 - 0.362 0.643
Weather 0.139 0.140 0.143 0.150 0.144 - 0.164 0.135 0.132 0.143 0.147 0.152 0.151 0.174 0.143 0.144 0.174 0.214 0.217 0.185 0.217 0.217
Agg. Relative Score 0.645 0.662 0.667 0.678 0.687 0.804 1.097 0.696 0.720 0.684 0.733 0.842 0.639 0.757 0.672 0.681 0.728 0.838 0.793 0.761 1.000 1.152
Avg. Rank 8.333 9.407 9.889 11.296 11.185 15.352 18.148 11.778 11.259 7.037 8.333 11.407 7.111 12.333 9.037 8.741 10.093 10.852 10.444 12.778 18.667 19.519
@@
Australian Electricity 1.333 1.319 1.399 1.114 1.310 1.186 2.158 1.635 1.258 1.009 0.871 1.473 0.997 0.810 1.278 0.794 0.828 1.161 1.427 2.391 0.897 1.393 1.253 2.362
Car Parts 0.906 0.899 0.887 0.891 0.881 - 2.657 0.816 1.735 1.542 0.803 0.798 0.817 0.799 0.879 0.803 0.803 0.891 1.157 1.185 1.229 - 1.201 -
CIF 2016 0.986 0.981 0.989 1.051 1.046 1.384 3.588 2.235 1.197 1.160 1.537 1.363 1.309 1.553 1.145 1.389 1.440 0.960 0.907 0.957 1.002 1.006 1.289 1.263
Covid Deaths 42.550 42.687 42.670 43.621 48.215 32.143 91.515 78.456 33.062 33.108 36.465 38.203 102.457 30.635 40.418 31.771 31.730 75.909 33.595 38.114 45.407 31.705 46.912 46.912
Dominick 0.818 0.816 0.819 0.833 0.820 - 3.274 1.250 0.879 0.845 0.867 0.851 0.812 0.800 0.880 0.782 0.782 1.813 0.891 0.885 1.016 - 0.871 0.871
ERCOT Load 0.617 0.550 0.573 0.588 0.561 1.319 3.975 0.834 0.583 0.667 0.553 1.197 0.780 0.690 0.651 0.615 0.648 0.558 1.308 2.826 1.306 1.284 0.761 4.234
ETT (15 Min.) 0.741 0.739 0.710 0.792 0.796 1.042 1.138 0.967 0.981 0.753 0.652 0.874 1.339 0.962 0.724 0.643 0.659 0.574 0.673 1.183 0.583 0.879 1.169 1.164
ETT (Hourly) 0.735 0.789 0.789 0.797 0.768 1.232 1.833 1.002 0.902 0.845 0.729 0.814 1.509 0.875 0.695 0.811 0.782 0.768 0.850 1.139 0.900 0.977 0.932 1.651
Exchange Rate 2.375 2.433 2.252 2.030 2.335 1.743 7.583 3.087 1.507 1.909 1.540 1.615 3.105 2.361 1.459 2.041 2.149 2.709 1.749 1.643 1.648 1.882 1.740 1.874
FRED-MD 0.500 0.486 0.496 0.483 0.468 0.513 2.621 2.283 0.607 0.593 0.745 0.621 0.849 0.929 0.713 0.696 0.635 0.693 0.492 0.544 0.566 0.473 1.101 0.622
Hospital 0.810 0.810 0.815 0.817 0.831 0.861 1.775 0.939 0.821 0.826 0.859 0.804 0.857 0.799 0.940 0.781 0.760 0.793 0.748 0.760 0.761 0.820 0.921 0.968
M1 (Monthly) 1.090 1.117 1.169 1.174 1.182 1.415 2.172 1.875 1.272 1.238 1.208 1.122 1.266 1.326 1.369 1.333 1.236 1.198 1.023 1.072 1.099 1.153 1.314 1.468
M1 (Quarterly) 1.713 1.739 1.764 1.785 1.785 1.802 9.931 3.036 1.896 1.840 1.920 1.741 1.904 2.144 1.943 2.061 2.043 1.958 1.602 1.710 1.683 1.770 2.078 1.952
M1 (Yearly) 4.301 4.624 4.659 4.958 4.751 4.077 23.089 7.149 4.623 4.708 4.042 3.685 4.727 4.316 11.565 5.568 6.212 3.675 3.571 4.110 3.697 3.870 4.894 4.894
M3 (Monthly) 0.857 0.868 0.885 0.900 0.930 0.996 2.240 1.846 0.946 0.924 1.225 0.943 0.950 0.916 1.161 0.899 0.883 0.950 0.827 0.869 0.861 0.933 1.146 1.175
M3 (Quarterly) 1.181 1.199 1.256 1.289 1.241 1.450 10.176 2.886 1.428 1.429 1.264 1.209 1.257 1.160 1.572 1.202 1.147 1.448 1.135 1.125 1.130 1.419 1.425 1.464
M3 (Yearly) 3.106 3.209 3.276 3.385 3.158 3.140 18.728 5.114 3.661 3.822 2.949 2.827 3.026 2.860 3.435 3.432 3.547 3.418 2.703 2.696 2.613 3.165 3.172 3.172
M4 (Quarterly) 1.216 1.231 1.246 1.271 1.312 - 6.927 2.663 1.286 1.259 1.150 1.254 1.241 1.248 1.229 1.157 1.129 1.215 1.145 1.188 1.193 1.276 1.602 1.477
M4 (Yearly) 3.606 3.678 3.651 3.743 3.933 - - 5.866 3.599 4.175 3.072 3.178 3.221 3.119 3.295 - - 3.374 3.013 3.374 3.124 3.730 3.974 3.974
M5 0.944 0.939 0.940 0.944 0.969 - 1.530 0.965 1.442 0.929 0.919 0.956 0.959 0.909 1.027 0.917 0.917 0.935 1.096 1.101 1.100 1.057 1.399 1.399
NN5 (Daily) 0.573 0.585 0.615 0.642 0.601 0.953 1.375 0.992 0.698 0.625 0.575 0.585 0.585 0.556 0.604 0.571 0.571 0.720 1.052 1.039 1.073 1.214 1.292 1.292
NN5 (Weekly) 0.940 0.938 0.944 0.947 0.963 0.968 1.349 1.141 0.980 1.009 0.877 0.920 1.034 0.896 0.966 0.919 1.014 1.268 0.974 0.978 0.984 0.995 1.063 1.063
Tourism (Monthly) 1.761 1.828 1.900 1.950 1.783 2.139 4.348 3.030 2.039 1.910 1.572 1.529 1.629 1.686 1.551 1.514 1.486 1.573 1.441 1.497 1.680 1.573 1.631 3.591
Tourism (Quarterly) 1.677 1.717 1.730 1.829 1.828 1.916 5.595 3.695 2.722 2.281 1.723 1.586 1.769 1.729 1.690 1.585 1.618 1.750 1.501 1.590 1.658 1.661 1.699 3.633
Tourism (Yearly) 3.755 3.900 3.901 4.048 3.862 3.309 12.093 3.755 3.047 3.301 3.138 3.702 4.130 3.047 3.406 3.448 3.564 - 3.276 3.138 3.078 4.043 3.552 3.552
Traffic 0.804 0.828 0.837 0.850 0.818 0.973 1.909 0.829 0.726 0.759 0.790 0.737 0.797 0.880 0.821 0.927 0.968 0.787 - 1.685 1.794 - 1.077 2.052
Weather 0.822 0.824 0.836 0.853 0.858 - 2.003 1.001 0.831 0.807 0.860 0.911 0.945 0.913 0.997 0.910 0.888 0.972 0.933 1.079 0.991 0.907 1.004 1.004
Agg. Relative Score 0.823 0.832 0.841 0.850 0.852 0.962 2.450 1.291 0.907 0.876 0.810 0.843 0.951 0.847 0.894 0.830 0.835 0.895 0.838 0.953 0.875 0.908 1.000 1.188
Avg. Rank 8.481 9.296 10.593 12.037 11.630 16.593 23.204 19.667 13.037 12.444 8.222 9.111 14.074 9.778 12.704 9.463 9.648 12.111 8.204 10.704 9.593 13.444 16.778 19.185